# Unsupervised Density Neural Representation for CT Metal Artifact Reduction

### v1

@@ IV-F3 Influence of Discretization of Energy Spectrum
The energy of a polychromatic X-ray physically covers a continuous range. However, due to limitations of physical instruments and numerical methods, measuring or estimating energy spectra are often based on discrete forms [42, 48, 49]. Our forward model $\mathcal{T}$ also uses the discrete energy spectrum (Eq. (11)). Here, we discuss how this discretization affects the MAR performance. We set the energy range to [20, 120] KeV and then uniformly sample $N=\{1,5,10,40,81,101\}$ energy levels from the range. Figure 12 shows the performance curve of our Diner over the discrete energy spectrum. From $N=5$ to $N=101$, the model performance remains excellent and robust (the change of the PSNR is less than 2 dB). When $N$ is set to 1, our Diner fails to produce good results (PSNR is only about 27 dB). This is because the forward model $\mathcal{T}$ degrades to a linear integral model (Eq. (3)) when $N=1$, in which the energy-induced BHE cannot be handled. The qualitative results are shown in Figure 13. Visually, the MAR reconstruction is satisfactory in both local details and global structures when $N=40$, while it contains several metal artifacts when $N=1$. In summary, the MAR performance of our Diner is not affected by the discretization of the energy spectrum.
Figure 12: Performance curve of our Diner with different discrete energy spectra $N=\{101,81,40,10,5,1\}$ on the DeepLesion dataset [43]. Note that $N=1$ means the degradation of the proposed forward model $\mathcal{T}$ towards the linear integral model (Eq. (2)).
Figure 13: Qualitative results of our Diner with two discrete energy spectra $N=\{1,40\}$ on a sample of the DeepLesion dataset [43]. Note $N=1$ means the degradation of the proposed forward model $\mathcal{T}$ towards the linear integral model (Eq. (2)). The red regions denote metals.
@@ IV-G Failure Cases
Our Diner proposes a novel nonlinear forward model $\mathcal{T}$ (Eq. (12)) to account for the complex BHE. However, there is another challenge when imaging objects containing ultra-high absorption metals, such as gold. As shown in Figure 14a, the X-rays hitting the ultra-high absorption metals are almost absorbed, resulting in the loss of valid measurement signals at the corresponding locations of the detectors. This phenomenon is called the photon starvation effect. Figure 14b shows a simulated gold-corrupted CT measurement, where the red regions represent the invalid signals and are background noise. However, our Diner considers the entire measurements as valid signals and uses them to optimize the INR network, causing the gradient error propagation. We show the MAR reconstructions by the SOTA-supervised ACDNet [9] and our Diner for the gold-corrupted CT measurement in Figure 14 (bottom row). Visually, the results of both techniques contain severe shadow artifacts and are unacceptable image quality. In summary, our Diner fails to handle the photon starvation effect caused by the ultra-high absorption metals. Nevertheless, we would like to emphasize that the common medical metal implants (e.g., titanium, chromium, and 304 stainless steel) do not absorb X-rays as well as gold. Therefore, the proposed approach works for most clinical scenarios.

### v2

@@ IV-F3 Influence of Discretization of Energy Spectrum
The energy of a polychromatic X-ray physically covers a continuous range. However, due to limitations of physical instruments and numerical methods, measuring or estimating energy spectra are often based on discrete forms [42, 45, 46]. Our forward model $\mathcal{T}$ also uses the discrete energy spectrum (Eq. 11). Here, we discuss how this discretization affects the MAR performance. We set the energy range to [20, 120] KeV and then uniformly sample $N=\{1,5,10,40,81,101\}$ energy levels from the range. Figure 12 shows the performance curve of our Diner over the discrete energy spectrum. From $N=5$ to $N=101$, the model performance remains excellent and robust (the change of the PSNR is less than 2 dB). When $N$ is set to 1, our Diner fails to produce good results (PSNR is only about 27 dB). This is because the forward model $\mathcal{T}$ degrades to a linear integral model (Eq. 3) when $N=1$, in which the energy-induced BHE cannot be handled. The qualitative results are shown in Figure 13. Visually, the MAR reconstruction is satisfactory in both local details and global structures when $N=40$, while it contains several metal artifacts when $N=1$. In summary, the MAR performance of our Diner is not affected by the discretization of the energy spectrum.
Figure 12: Performance curve of our Diner with different discrete energy spectra $N=\{101,81,40,10,5,1\}$ on the DeepLesion dataset [43]. Note that $N=1$ means the degradation of the proposed forward model $\mathcal{T}$ towards the linear integral model (Eq. 2).
Figure 13: Qualitative results of our Diner with two discrete energy spectra $N=\{1,40\}$ on a sample of the DeepLesion dataset [43]. Note $N=1$ means the degradation of the proposed forward model $\mathcal{T}$ towards the linear integral model (Eq. 2). The red regions denote metals.
@@ IV-G Failure Cases
Our Diner proposes a novel nonlinear forward model $\mathcal{T}$ (Eq. 12) to account for the complex BHE. However, there is another challenge when imaging objects containing ultra-high absorption metals, such as gold. As shown in Figure 14a, the X-rays hitting the ultra-high absorption metals are almost absorbed, resulting in the loss of valid measurement signals at the corresponding locations of the detectors. This phenomenon is called the photon starvation effect. Figure 14b shows a simulated gold-corrupted CT measurement, where the red regions represent the invalid signals and are background noise. However, our Diner considers the entire measurements as valid signals and uses them to optimize the INR network, causing the gradient error propagation. We show the MAR reconstructions by the SOTA-supervised ACDNet [8] and our Diner for the gold-corrupted CT measurement in Figure 14 (bottom row). Visually, the results of both techniques contain severe shadow artifacts and are unacceptable image quality. In summary, our Diner fails to handle the photon starvation effect caused by the ultra-high absorption metals. Nevertheless, we would like to emphasize that the common medical metal implants (e.g., titanium, chromium, and 304 stainless steel) do not absorb X-rays as well as gold. Therefore, the proposed approach works for most clinical scenarios.